• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kuss

    Log entry time 09:29:05 on March 7,2000

    Entry number 35772

    keyword=target density change

    Lowered the target density from 155psi to 66.4psi. According to the diagram in the counting house this is a drop from 68mg/cm3 to 34mg/cm3, exactly 50% (this was pure luck).

    Both fans are running at 25%, JT valve settings are 25% (JT1) and 30% (JT3). Both heaters together deposit 200W heat load, which is the equivalent of a 160µA beam current, at this density (1.2W/µA).

    Target is ready.